# Arens extensions of disjointness preserving multilinear operators on Riesz spaces

### v1

@@ Abstract
Let $E_{1},\ldots,E_{m}$ be (non necessarily Archimedean) Riesz spaces, let $F$ be an Archimedean Riesz space and let $A\colon E_{1}\times\cdots\times E_{m}\to F$ be an order bounded disjointness preserving $m$-linear operator. We prove that all Arens extensions of $A$ are disjointness preserving if either $A$ has finite rank or the spaces are Banach lattices and $F^{*}$ has a Schauder basis consisting of disjointness preserving functionals.
@@ 1 Introduction
The study of disjointness preserving linear operators between Riesz spaces (vector lattices) started with the seminal work of Abramovich, Veskler and Koldunov [4] in 1979. Next, several authors investigated different aspects of such class of operators, see, e.g., [2, 3, 9, 12, 13, 19]. Multilinear disjointness preserving operators were considered later by Kusraev and Tabuev [15] in 2004, and then studied by Kusraev and Kusraeva in [14], by Toumi in [21], and, more recently, by Kusraeva in [16].
The main purpose of this work is to establish conditions so that Arens extensions of disjointness preserving multilinear operators between Riesz spaces are also disjointness preserving. The corresponding problems for Riesz multimorphisms, for almost Dunford-Pettis multilinear operators and for order continuous multilinear operators were addressed by the first two authors, respectively, in [6, 7, 8]. In our way to solve the aforementioned problem, we need first to study the linear case and to establish basic facts about disjointness preserving multilinear operators. This is done in Section 2, where, in particular, we: (i) Give conditions so that the second adjoint of a disjointess preserving linear operator $T\colon E\to F$ between Riesz spaces is also disjointness preserving. (ii) Prove that if $E_{1},\dots,E_{m}$ are Riesz spaces and $A\colon E_{1}\times\cdots\times E_{m}\to\mathbb{R}$ is a positive disjointess $m$-linear form, then there exist disjointness preserving positive linear functionals $\varphi_{1}\in E_{1}^{\sim},\ldots,\varphi_{m}\in E_{m}^{\sim}$ such that $A(x_{1},\ldots,x_{m})=\varphi_{1}(x_{1})\cdots\varphi_{m}(x_{m})$ for all $x_{1}\in E_{1},\ldots,x_{m}\in E_{m}$. The main results, on disjointness preservation of Arens extensions of disjointness preserving multilinear operators, are proved in Section 3. It is important to mention that, in most of our results, the Riesz spaces on the domain of linear/multilinear operators are not supposed to be Archimedean as they were in [9, 14, 16, 19]. The case of operators between Banach lattices is also considered.
We refer the reader to [5, 18] for background on Riesz spaces and Banach lattices. Now, we present some terminology considering multilinear maps between Riesz spaces. Given Riesz spaces $E_{1},\ldots,E_{m},F$, the space of all order bounded (respectively, regular) $m$-linear operators from $E_{1}\times\cdots\times E_{m}$ to $F$ is denoted by ${\cal L}_{b}(E_{1},\ldots,E_{m};F)$ (respectively, ${\cal L}_{r}(E_{1},\ldots,E_{m};F)$). For precise definitions, see [14, 17]. When $F=\mathbb{R}$, we write ${\cal L}_{b}(E_{1},\ldots,E_{m})$ (respectively, ${\cal L}_{r}(E_{1},\ldots,E_{m})$). The order relation in these spaces is given by
@@
As usual, in the linear case $m=1$, the order dual of a Riesz space $E$ is denoted by $E^{\sim}$. The set of all permutations of $\{1,\ldots,m\}$ is denoted by $S_{m}$.
@@ 2 Disjointness preserving linear and multilinear operators
Recall that the (order) adjoint of an order bounded linear operator $T\colon E\to F$ between two Riesz spaces is the restriction of the adjoint of $T$ to $F^{\sim}$, denoted by $T^{\prime}\colon F^{\sim}\to E^{\sim}$. Of course, $T{{}^{\prime\prime}}:=(T^{\prime})^{\prime}$. As announced in the Introduction, we begin this section of preparatory results studying when the second adjoint $T^{\prime\prime}\colon E^{\sim\sim}\to F^{\sim\sim}$ of a disjointness preserving linear operator $T\colon E\to F$ between two Riesz spaces is disjointness preserving as well. Let us recall that a linear operator $T\colon E\to F$ between two Riesz spaces is said to be disjointness preserving if $|T(x)|\wedge|T(y)|=0$ for all $x,y\in E$ with $|x|\wedge|y|=0$.
@@
(1) By [5, Theorem 2.40] the modulus of $T$ exists and $|T|(|x|)=|T(|x|)|=|T(x)|$. Applying [5, Theorem 1.7] we get
@@
(3) Let $x,y\in E$ with $|x|\wedge|y|=0$. As $J_{E}\colon E\to E^{\sim\sim}$ is a Riesz hommorphism, we have $|J_{E}(x)|\wedge|J_{F}(y)|=0$, from which it follows that $|T^{\prime\prime}(J_{E}(x))|\wedge|T^{\prime\prime}(J_{E}(y))|=0$. Since $T^{\prime\prime}(J_{E}(z))=J_{F}(T(z))$ holds for every $z\in E$, we obtain that $|J_{F}(T(x))|\wedge|J_{F}(T(y))|=0$ in $F^{\sim\sim}$. Finally, since $F$ is a Riesz subspace of $F^{\sim\sim}$, $|T(x)|\wedge|T(y)|=0$ and we are done. ∎
@@
It is well known that a positive linear operator between Archimedean Riesz spaces is disjointness preserving if and only if it is a Riesz homomorphism (see, e.g., [19, p. 543]). Next we show that no condition on the underlying Riesz spaces is required for this to be true.
@@
A positive linear operator $T\colon E\to F$ between Riesz spaces is disjointness preserving if and only if $T$ is a Riesz homomorphism. In this case, $T^{\prime\prime}\colon E^{\sim\sim}\to F^{\sim\sim}$ is also a Riesz homomorphism, hence disjointness preserving.
@@
Assuming that $T$ preserves disjointness, $|T(x^{+})|\wedge|T(x^{-})|=0$ holds for every $x\in E$. So, by [5, Theorem 1.7(5)],
$0=|T(x^{+})|\wedge|T(x^{-})|=\dfrac{1}{2}\big{|}|T(x^{+})+T(x^{-})|-|T(x^{+})-T(x^{-})|\big{|}=\dfrac{1}{2}\big{|}T(|x|)-|T(x)|\big{|}.$
Hence, $T(|x|)=|T(x)|$ for every $x\in E$, from which it follows that $T$ is a Riesz homomorphism. The converse is obvious. Calling on [5, Theorem 2.19 and 2.20], its second adjoint $T^{\prime\prime}$ is a Riesz homomorphism as well. Note that [5, Theorem 2.19] and the implication of [5, Theorem 2.20] we are using do not require any assumption on the spaces. In particular, $T^{\prime\prime}$ preserves disjointness. ∎
The following result, which is a multilinear version of [5, Theorem 1.14], shall be helpful in our study about disjointness preserving multilinear operators.
Theorem 2.4.
Let $E_{1},\ldots,E_{m},F$ be Riesz spaces with $F$ Archimedean and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ be an $m$-linear operator. Suppose that, regardless of the positive vectors $x_{1}\in E_{1},\dots,x_{m}\in E_{m}$,
[…]
exists in $F$. Then the modulus of $A$ exists and $|A|(x_{1},\dots x_{m})$ coincides with the supremum above for all $x_{1}\in E_{1}^{+},\dots,x_{m}\in E_{m}^{+}.$
Proof.
We prove the bilinear case $m=2$. The general case follows analogously. By assumption, the map $B\colon E_{1}^{+}\times E_{2}^{+}\to F^{+}$ given by
$B(x_{1},x_{2})=\sup\left\{{|A(z_{1},z_{2})|}\,:\,{|z_{1}|\leq x_{1},\,|z_{2}|\leq x_{2}}\right\}$
is well defined. Reasoning as in the proof of [5, Theorem 1.14], $|A(z_{1},z_{2})|$ can be replace with $A(z_{1},z_{2})$ in the definition of $B$. Let us check now that $B$ is additive in each variable. Fix $y\in E_{2}^{+}$ and take $u,v\in E_{1}^{+}$. If $|u_{0}|\leq u$ and $|v_{0}|\leq v$, then $|u_{0}+v_{0}|\leq|u_{0}|+|v_{0}|\leq u+v$, which implies that
$A(u_{0},y)+A(v_{0},y)=A(u_{0}+v_{0},y)\leq B(u+v,y).$
A straightforward argument shows that $B(u,y)+B(v,y)\leq B(u+v,y)$. On the other hand, if $|w|\leq u+v$, there exits $u_{0},v_{0}\in E$ such that $|u_{0}|\leq u$, $|v_{0}|\leq v$ and $w=u_{0}+v_{0}$ (see [5, Theorem 1.13]). Hence,
$A(w,y)=A(u_{0},y)+A(v_{0},y)\leq B(u,y)+B(u,y),$
which gives that $B(u+v,w)\leq B(u,w)+B(v,w)$. Therefore $B(u+v,w)=B(u,w)+B(v,w)$. The same argument shows that $B$ is additive in the second variable as well. By [17, Theorem 2.3], $B$ can be extended to a positive bilinear map from $E_{1}\times E_{2}$ to $F$, which is also denote by $B$.
Now we prove that $B=\sup\{-A,A\}$. First, we observe that $-A\leq B$ and $A\leq B$ holds immediately. Let $T\colon E_{1}\times E_{2}\to F$ be a bilinear map such that $-A\leq T$ and $A\leq T$. It is clear that $T\geq 0$. Fix $x\in E_{1}^{+}$ and $y\in E_{2}^{+}$. Given $|u|\leq x$ and $|v|\leq y$, we have
$A(u,v)$ $=A(u^{+},v^{+})-A(u^{+},v^{-})-A(u^{-},v^{+})+A(u^{-},v^{-})$
$\leq T(u^{+},v^{+})+T(u^{+},v^{-})+T(u^{-},v^{+})+T(u^{-},v^{-})$
$=T(|u|,|v|)\leq T(x,y).$
Taking the supremum over all $|u|\leq x$ and $|v|\leq y$, we get $B(x,y)\leq T(x,y)$, which proves that $B=\sup\{-A,A\}=|A|$. ∎
From now on we focus on disjointness preserving multilinear operators. Given Riesz spaces $E_{1},\dots,E_{m},$ and $F$, an $m$-linear operator $A\colon E_{1}\times\cdots\times E_{m}\to F$ is said to preserve disjointness if, for all $j\in\{1,\ldots,m\}$ and $x_{i}\in E_{i}^{+},\,i\neq j$, the linear operator
$A_{j}:=A(x_{1},\ldots,x_{j-1},\cdot,x_{j+1},\ldots,x_{m})\colon E_{j}\longrightarrow F,\,A_{j}(x_{j})=A(x_{1}\ldots,x_{m}),$
preserves disjointness.
Our first result on disjointness preserving multilinear operators is a multilinear counterpart of [5, Theorem 2.40].
Proposition 2.5.
Let $E_{1},\ldots,E_{m},F$ be Riesz spaces with $F$ Archimedean and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ be an order bounded $m$-linear operator. If $A$ is disjointness preserving, then its modulus exists and
for all $x_{1}\in E_{1},\ldots,x_{m}\in E_{m}$. In particular, $|A|$ is disjointness preserving.
Proof.
Again, for simplicity, we prove the bilinear case $m=2$. We begin by noticing that, for all $a_{1}\in E_{1}^{+}$ and $a_{2}\in E_{2}^{+}$, the order bounded linear operators $A_{a_{1}}\colon E_{2}\longrightarrow F,A_{a_{1}}(x_{2})=A(a_{1},x_{2})$, and $A^{a_{2}}\colon E_{1}\longrightarrow F,A^{a_{2}}(x_{1})=A(x_{1},a_{2})$, preserve disjointness. So, by [5, Theorem 2.40] their modulus exist and
@@
for all $x_{1}\in E_{1}$ and $x_{2}\in E_{2}$. Given $x_{1}\in E_{1}^{+}$, $x_{2}\in E_{2}^{+}$, $|z_{1}|\leq x_{1}$ and $|z_{2}|\leq x_{2}$, from (1) it follows that
$|A(z_{1},z_{2})|$ $=|A(z_{1}^{+},z_{2})-A(z_{1}^{-},z_{2})|\leq|A(z_{1}^{+},z_{2})|+|A(z_{1}^{-},z_{2})|$
$=|A_{z_{1}^{+}}|(|z_{2}|)+|A_{z_{1}^{-}}|(|z_{2}|)\leq|A_{z_{1}^{+}}|(x_{2})+|A_{z_{1}^{-}}|(x_{2})$
$=|A(z_{1}^{+},x_{2})|+|A(z_{1}^{-},x_{2})|=|A(z_{1}^{+},x_{2})+A(z_{1}^{-},x_{2})|$
$=|A(|z_{1}|,x_{2})|=|A^{x_{2}}|(|z_{1}|)$
$\leq|A^{x_{2}}|(x_{1})=|A(x_{1},x_{2})|,$
which implies that
$|A(x_{1},x_{2})|\leq\sup\left\{{|A(z_{1},z_{2})|}\,:\,{|z_{i}|\leq x_{i},\,i=1,2}\right\}\leq|A(x_{1},x_{2})|,$
proving that $\sup\left\{{|A(z_{1},z_{2})|}\,:\,{|z_{i}|\leq x_{i},\,i=1,2}\right\}=|A(x_{1},x_{2})|$ in $F$ for all $x_{1}\in E_{1}^{+}$ and $x_{2}\in E_{2}^{+}$. By Theorem 2.4, the modulus of $A$ exists and
$|A|(x_{1},x_{2})=\sup\left\{{|A(z_{1},z_{2})|}\,:\,{|z_{i}|\leq x_{i},\,i=1,2}\right\}=|A(x_{1},x_{2})|$
for all $x_{1}\in E_{1}^{+}$ and $x_{2}\in E_{2}^{+}$. In particular, $|A|(|x_{1}|,|x_{2}|)=|A(|x_{1}|,|x_{2}|)|$ holds for all $x_{1}\in E_{1}$ and $x_{2}\in E_{2}$. Therefore, using $|A(x_{1},x_{2}^{+})|\wedge|A(x_{1},x_{2}^{-})|=0$ and (1),
$|A(x_{1},x_{2})|$ $=|A(x_{1},x_{2}^{+})-A(x_{1},x_{2}^{-})|=|A(x_{1},x_{2}^{+})-A(x_{1},x_{2}^{-})|=|A(x_{1},|x_{2}|)|=|A(|x_{1}|,|x_{2}|)|$
holds for all $x_{1}\in E_{1}$ and $x_{2}\in E_{2}$. Then, for $x_{1}\in E_{1}$ and $x_{2}\in E_{2}$,
@@
$=|A(x_{1}^{+},x_{2}^{+})|-|A(x_{1}^{+},x_{2}^{-})|-|A(x_{1}^{-},x_{2}^{+})|+|A(x_{1}^{-},x_{2}^{-})|.$
Since $A$ is disjointness preserving, $|A(x_{1}^{+},x_{2}^{+})|\wedge|A(x_{1}^{-},x_{2}^{+})|=0$, so, by [5, Theorem 1.7(3)], we get

### v2

@@ Abstract
Let $E_{1},\ldots,E_{m}$ be (non necessarily Archimedean) Riesz spaces, let $F$ be a Dede-kind complete Riesz space and let $A\colon E_{1}\times\cdots\times E_{m}\to F$ be a regular disjointness preserving $m$-linear operator. We prove that all Arens extensions of $A$ are disjointness preserving if either $A$ has finite rank or the spaces are Banach lattices and $F^{*}$ has a Schauder basis consisting of disjointness preserving functionals.
@@ 1 Introduction
The study of disjointness preserving linear operators between Riesz spaces (vector lattices) started with the seminal work of Abramovich, Veskler and Koldunov [4] in 1979. Next, several authors investigated different aspects of such class of operators, see, e.g., [2, 3, 9, 14, 15, 23]. Multilinear disjointness preserving operators were considered later by Kusraev and Tabuev [18] in 2004, and then studied by Bu, Buskes and Kusraev in [11], by Kusraev and Kusraeva in [16], by Toumi in [25], and, more recently, by Kusraeva in [19, 20].
The main purpose of this work is to establish conditions so that Arens extensions of disjointness preserving multilinear operators between Riesz spaces are also disjointness preserving. The corresponding problems for Riesz multimorphisms, for almost Dunford-Pettis multilinear operators and for order continuous multilinear operators were addressed by the first two authors, respectively, in [6, 7, 8]. In our way to solve the aforementioned problem, we need first to study the linear case and to establish basic facts about disjointness preserving multilinear operators. This is done in Section 2, where, in particular, we give conditions so that the second adjoint of a disjointess preserving linear operator $T\colon E\to F$ between Riesz spaces is also disjointness preserving. Our main results, announced in the Abstract, on Arens extensions of disjointness preserving multilinear operators between Riesz spaces/Banach lattices are proved in Section 3. It is important to mention that, in most of our results, the Riesz spaces on the domain of linear/multilinear operators are not supposed to be Archimedean as they were in [9, 16, 19, 23].
We refer the reader to [5, 22] for background on Riesz spaces and Banach lattices. Now, we present some terminology considering multilinear maps between Riesz spaces. Given Riesz spaces $E_{1},\ldots,E_{m},F$, the space of all regular $m$-linear operators from $E_{1}\times\cdots\times E_{m}$ to $F$ is denoted by ${\cal L}_{r}(E_{1},\ldots,E_{m};F)$. For precise definitions, see [21]. When $F=\mathbb{R}$, we write ${\cal L}_{r}(E_{1},\ldots,E_{m})$. The order relation in this space is given by
@@
The positive cone of a Riesz space $E$ is denoted by $E^{+}$. As usual, in the linear case $m=1$, the order dual of a Riesz space $E$ is denoted by $E^{\sim}$ and the topological dual of a Banach lattice $E$ by $E^{*}$. The symbol $x\perp y$ means that the vectors $x$ and $y$ of a Riesz space are disjoint. The set of all permutations of $\{1,\ldots,m\}$ is denoted by $S_{m}$.
@@ 2 Disjointness preserving linear and multilinear operators
Recall that the (order) adjoint of an order bounded linear operator $T\colon E\to F$ between two Riesz spaces is the restriction of the adjoint of $T$ to $F^{\sim}$, denoted by $T^{\prime}\colon F^{\sim}\to E^{\sim}$. Of course, $T{{}^{\prime\prime}}:=(T^{\prime})^{\prime}$. As announced in the Introduction, we begin this section of preparatory results studying when the second adjoint $T^{\prime\prime}\colon E^{\sim\sim}\to F^{\sim\sim}$ of a disjointness preserving linear operator $T\colon E\to F$ between two Riesz spaces is disjointness preserving as well. Let us recall that a linear operator $T\colon E\to F$ between two Riesz spaces is said to be disjointness preserving if $T(x)\perp T(y)$ in $F$ whenever $x\perp y$ in $E$.
@@
(1) By [5, Theorem 2.40] the modulus of $T$ exists and $|T|(|x|)=|T(|x|)|=|T(x)|$ for every $x\in E$. Applying [5, Theorem 1.7] we get
@@
(3) Let $x$ and $y$ be disjoint vectors in $E$. As $J_{E}\colon E\to E^{\sim\sim}$ is a Riesz homomorphism, we have $J_{E}(x)\perp J_{F}(y)$, from which it follows that
$J_{F}(T(x))=T^{\prime\prime}(J_{E}(x))\perp T^{\prime\prime}(J_{E}(y))=J_{F}(T(y))\mbox{ in }F^{\sim\sim}$
because $T^{\prime\prime}$ preserves disjointness. Using now that $(J_{F})^{-1}\colon J_{F}(F)\to F$ is a Riesz homomorphism we get $T(x)\perp T(y)$. ∎
@@
Remark 2.3.
The situation is quite simpler for positive operators. The following are equivalent for a positive linear operator $T\colon E\to F$ between Riesz spaces.
(a) $T$ is disjointness preserving.
(b) $T$ is a Riesz homomorphism.
(c) $T^{\prime\prime}\colon E^{\sim\sim}\to F^{\sim\sim}$ is disjointness preserving.
(d) $T^{\prime\prime}\colon E^{\sim\sim}\to F^{\sim\sim}$ is a Riesz homomorphism.
Indeed, (a)$\Leftrightarrow$(b) is well known (and it is easy to be checked directly); so (c)$\Leftrightarrow$(d) also holds because $T^{\prime\prime}\geq 0$. (a)$\Rightarrow$(d) follows from a combination of [5, Theorem 2.19], which imposes no conditions on $E$ and $F$, with the implication of [5, Theorem 2.20] that does not require any assumption on the spaces. For (d)$\Rightarrow$(a), use that $E$ and $F$ are (Riesz isomorphic to) sublattices of $E^{\sim\sim}$ and $F^{\sim\sim}$, or, alternatively, that $T=(J_{F})^{-1}\circ T^{\sim\sim}\circ J_{E}$, $(J_{F})^{-1}\colon J_{F}(F)\to F$.
From now on we focus on disjointness preserving multilinear operators, a topic for which we gave several references in the Introduction. Given Riesz spaces $E_{1},\dots,E_{m},$ and $F$, an $m$-linear operator $A\colon E_{1}\times\cdots\times E_{m}\to F$ is said to preserve disjointness if $A$ preserves disjointness separately for fixed positive vectors, that is, for all $j\in\{1,\ldots,m\}$ and $x_{i}\in E_{i}^{+},\,i\neq j$, the linear operator
$x_{j}\in E_{j}\longrightarrow A(x_{1}\ldots,x_{m})\in F$
preserves disjointness. It is not difficult to check that this is equivalent to say that $A$ preserves disjointness separately, meaning that the above holds for all $j\in\{1,\ldots,m\}$ and $x_{i}\in E_{i},i\neq j$.
It is stated in [16, Theorem 3] that, for every order bounded disjointness preserving $m$-linear operator $A\colon E_{1}\times\cdots\times E_{m}\to F$ between Archimedean Riesz spaces, $|A|$ exists and satisfies
for all $x_{1}\in E_{1},\ldots,x_{m}\in E_{m}$. Since [16, Theorem 3] is just a restatement of a result in the difficult-to-read reference [18] (it is in Russian), we prove next a similar version for regular operators that shall fit our purposes.
Recall that an $m$-linear operator $A\colon E_{1}\times\cdots\times E_{m}\to F$ between Riesz spaces is a Riesz multimorphism (or an $m$-morphism) if $|A(x_{1},\ldots,x_{m})|=A(|x_{1}|,\ldots,|x_{m}|)$ for all $x_{1}\in E_{1},\ldots,x_{m}\in E_{m}$. Or, equivalently, if $A$ is a Riesz homomorphism separately for fixed positive vectors. As in the linear case, a positive multilinear operator preserves disjointness if and only if it is a Riesz multimorphism.
A positive partition of a positive vector $x$ of a Riesz space $E$ is a finite sequence of elements of $E^{+}$ whose sum equals $x$. The set of all positive partitions of $x$ shall be denoted by $\prod_{x}$.
@@
Let $E_{1},\ldots,E_{m},F$ be Riesz spaces with $F$ Dedekind complete and let $A\colon E_{1}\times\cdots\times E_{m}\longrightarrow F$ be a regular $m$-linear operator. If $A$ is disjointness preserving, then
for all $x_{1}\in E_{1},\ldots,x_{m}\in E_{m}$. Moreover,
$A^{+}(x_{1},\ldots,x_{m})=(A(x_{1},\ldots,x_{m}))^{+}\mbox{ and~{}}A^{-}(x_{1},\ldots,x_{m})=(A(x_{1},\ldots,x_{m}))^{-}$
for all $x_{1}\in E_{1}^{+},\ldots,x_{m}\in E_{m}^{+}$. In particular, $|A|$, $A^{+}$ and $A^{-}$ are Riesz multimorphisms.
@@
For simplicity, we prove the bilinear case $m=2$. Since $A$ is regular and $F$ is Dedekind complete, the modulus $|A|$ of $A$ exists and satisfies
$|A|(x_{1},x_{2})=\sup\left\{{\sum_{i_{1}=1}^{n_{1}}\sum_{i_{2}=1}^{n_{2}}|A(u_{i_{1}},v_{i_{2}})|}\,:\,{(u_{i_{1}})_{i_{1}=1}^{n_{1}}\in{\textstyle\prod_{x_{1}}},(v_{i_{2}})_{i_{2}=1}^{n_{2}}\in{\textstyle\prod_{x_{2}}}}\right\}$
for all $x_{1}\in E_{1}^{+},x_{2}\in E_{2}^{+}$ (see [21, Proposition 2.14]). Now, for all $a_{1}\in E_{1}^{+}$ and $a_{2}\in E_{2}^{+}$, the regular linear operators $A_{a_{1}}\colon E_{2}\longrightarrow F$ given by $A_{a_{1}}(x_{2})=A(a_{1},x_{2})$ and $A^{a_{2}}\colon E_{1}\longrightarrow F$ given by $A^{a_{2}}(x_{1})=A(x_{1},a_{2})$ preserve disjointness by assumption. By [5, Theorem 2.40] their moduli exist and
@@
for all $x_{1}\in E_{1}$ and $x_{2}\in E_{2}$. Given $x_{1}\in E_{1}^{+}$, $x_{2}\in E_{2}^{+}$, $(u_{i})_{i=1}^{n_{1}}\in\prod_{x_{1}}$ and $(v_{j})_{j=1}^{n_{2}}\in\prod_{x_{2}}$, from (1) it follows that
$\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}|A(u_{i},v_{j})|$ $=\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}|A_{u_{i}}|(v_{j})=\sum_{i=1}^{n_{1}}|A_{u_{i}}|\left(\sum_{j=1}^{n_{2}}v_{j}\right)$
$=\sum_{i=1}^{n_{1}}|A_{u_{i}}|(x_{2})=\sum_{i=1}^{n_{1}}|A(u_{i},x_{2})|$
$=\sum_{i=1}^{n_{1}}|A^{x_{2}}|(u_{i})=|A^{x_{2}}|\left(\sum_{i=1}^{n_{1}}u_{i}\right)$
$=|A^{x_{2}}|(x_{1})=|A(x_{1},x_{2})|.$
This implies that $|A(x_{1},x_{2})|=|A|(x_{1},x_{2})$ for all $x_{1}\in E_{1}^{+}$ and $x_{2}\in E_{2}^{+}$. In particular, $|A|(|x_{1}|,|x_{2}|)=|A(|x_{1}|,|x_{2}|)|$ holds for all $x_{1}\in E_{1}$ and $x_{2}\in E_{2}$. Therefore, using that $A(x_{1},x_{2}^{+})\perp A(x_{1},x_{2}^{-})$ and (1),
$|A(x_{1},x_{2})|$ $=|A(x_{1},x_{2}^{+})-A(x_{1},x_{2}^{-})|$
$=|A(x_{1},x_{2}^{+})+A(x_{1},x_{2}^{-})|=|A(x_{1},|x_{2}|)|=|A(|x_{1}|,|x_{2}|)|$
holds for all $x_{1}\in E_{1}$ and $x_{2}\in E_{2}$. Then,
@@
$=|A(x_{1}^{+},x_{2}^{+})|-|A(x_{1}^{+},x_{2}^{-})|-|A(x_{1}^{-},x_{2}^{+})|+|A(x_{1}^{-},x_{2}^{-})|$
for all for $x_{1}\in E_{1}$ and $x_{2}\in E_{2}$. Since $A$ is disjointness preserving, $A(x_{1}^{+},x_{2}^{+})\perp A(x_{1}^{-},x_{2}^{+})$, so, by [5, Theorem 1.7(3)] we get